ABC Company purchased $45,051 of equipment 4 years ago. The equipment is 7-year MACRS property. The firm is selling this equipment today for $4,789. What is the After-tax Salvage Value if the tax rate is 29%?
Answer #1

Book value as on date of sale=Cost-Accumulated depreciation

=45051(1-0.1429-0.2449-0.1749-0.1249)

=$14073.9324

Hence loss on sale=(14073.9324-4789)=$9284.9324

Hence after-tax salvage value=Sale proceeds+(Tax rate*Loss on sales)

=4789+(9284.9324*0.29)

which is equal to

=$7481.63(Approx).